OpenStudy (ayanaortega):

help please What is the discriminant of the given equation? X^2-2x+4=0 A. -12 B. 0 C. 20 D. 12

OpenStudy (ayanaortega):

@phi

OpenStudy (phi):

do your notes tell you what the discriminant is ?

OpenStudy (ayanaortega):

yes , here is what it said. The discriminant is the name given to the expression that appears under the square root (radical) sign in the quadratic formula. The discriminant tells you about the "nature" of the roots of a quadratic equation given that a, b and c are rational numbers.

OpenStudy (ayanaortega):

@phi

OpenStudy (anonymous):

do you know what is the formula of discriminant ?

OpenStudy (anonymous):

D = b^2 -4ac yes ?

OpenStudy (ayanaortega):

yes, Discriminant of a Quadratic. The number D = b2 – 4ac determined from the coefficients of the equation ax2 + bx + c = 0. The discriminant reveals what type of roots the equation has. Note: b2 – 4ac comes from the quadratic formula.

OpenStudy (anonymous):

yes right so what you need to know than about it ?

OpenStudy (ayanaortega):

how to find the answer

OpenStudy (anonymous):

so you know that D = b^2 -4ac in this case of your exercise a= ? b=? c= ?

OpenStudy (anonymous):

ax^2 +bx + c = 0 how is this in case of this posted quadratic ?

OpenStudy (ayanaortega):

0

OpenStudy (anonymous):

x^2 -2x +4 = 0 so can you tel me now how many is a,b and c equal ?

OpenStudy (anonymous):

so a= 1 bc. there is 1x^2 yes ?

OpenStudy (anonymous):

please collaborate

OpenStudy (ayanaortega):

a=1 b=2 c=4

OpenStudy (anonymous):

b not is 2

OpenStudy (anonymous):

bc than you write just 2 so tis mean +2 but there is -2 yes ?

OpenStudy (anonymous):

so now than you know these values of a,b and c so you can calculi easy the value of discriminant yes ?

OpenStudy (ayanaortega):

so it would be c right

OpenStudy (mrnood):

you now have the information needed a=1 b=-2 c=4 all you need to do is put those numbers into the expression D = b^2-4ac when you have worked out that value that is the answer